ConcreteT). MOE 29.8 33.2 33.7 35.3 35.5 36.1 36.2 36.3 37.5 37.7 38.7 38.8 39.6 41.0 Strength 5.9 7.2 7.3 6.3 8.1 6.8 7.0 7Comment on any interesting features. (Select all that apply.) I There is a small amount of variation in the data. There is aMS P Analysis of Variance SOURCE DE SS Regression 1 52.870 Error 25 18.736 Total 26 71.605 F 70.55 0.000 52.870 0.749 y = PreValues of modulus of elasticity (MOE, the ratio of stress, i.e., force per unit area, to strain, i.e., deformation per unit length, in GPa) and flexural strength (a measure of the ability to resist failure in bending, in MPa) were determined for a sample of concrete beams of a certain type, resulting in the following data (read from a graph in the article "Effects of Aggregates and Microfillers on the Flexural Properties of Concrete"†).
Answer #1

Ans a ) the stem and leaf display is

Stem-and-Leaf Display: MOE

Stem-and-leaf of MOE N = 25
Leaf Unit = 1.0

1 2
1 2 9
3 3 33
12 3 566677889
(4) 4 1223
9 4 5689
5 5 1
4 5
4 6 2
3 6 9
2 7
2 7 9
1 8 0

option A and E are true

b ) option A is true

c ) y = 3.2925+0.10748 x

for x= 37

predicted strengt = 3.2925+0.10748*37 =7.27

for x = 100

y = 3.2925+0.10748*100 = 14.04

option D is true

d )

SSE 18.736
SST 71.605
r^2 0.738

option b is true